OpenStudy (anonymous):

If f(x) = |x + 7|, find f(–12). A. –19 B. 5 C. –5 D. 19

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

f(x) = |x + 7| plugin x = -12 : f(-12) = |-12+7| = ?

OpenStudy (anonymous):

Plug -12 in for x f(-12) = |-12 + 7| f(-12) = |-5| f(-12) = 5 Option B

OpenStudy (anonymous):

So for all problems like this do you plug them in?

ganeshie8 (ganeshie8):

Yep